OpenStudy (anonymous):

What are the zeroes of f(x) = x^2 + 5x + 6? x = -2, -3 x = 2, 3 x = -2, 3 x = 2, -3

OpenStudy (kaederfds):

\[x^{2} + 5x + 6 = (x+3)(x+2)\] so x=-2,-3

OpenStudy (anonymous):

Thank you!!

OpenStudy (kaederfds):

you're welcome :)